OpenStudy (anonymous):

i realllly need help School Helene’s marching band needs money to travel to a competition. Band members have raised $560. They need to raise a total of $1680. Write and solve an equation to find how much more they need. Show that your answer is reasonable

OpenStudy (anonymous):

Do you have an answer yet?

OpenStudy (anonymous):

560+x=1680 -560 -560 x=$1120

OpenStudy (anonymous):

i got the same answer but i wasn't sure , thank you